Hypothesis test on population mean; 𝛼 = 0.05; one-tailed test (right); 𝑛 = 37; 𝑠 = 5


1
Expert's answer
2022-06-14T11:44:30-0400

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=36" and the critical value for a right-tailed test is "t_c =1.6882985."

The rejection region for this right-tailed test is "R = \{t:t>1.688298\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}"


If "t>1.688298=t_c," the null hypothesis is rejected.